OpenStudy (anonymous):

if a=-2 and b= -2, what is the value of (a^3-1) /(b-1)

OpenStudy (luigi0210):

\[\frac{ a^3-1 }{ b-1 }\] Is this your equation?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

well plug in 2 and solve ^_^

OpenStudy (anonymous):

Unfortunately we cannot give out answers, so where do you run into a problem stephanie?

OpenStudy (nurali):

\[\frac{ (-2)^3-1 }{ -2-1 }\]

OpenStudy (luigi0210):

since it's a a negative and cubed.. it'll be a negative

OpenStudy (nurali):

\[\frac{ -8-1 }{ -3 }\]

OpenStudy (luigi0210):

Like Nurali has shown :l

OpenStudy (nurali):

\[\frac{ 9 }{ 3 }\]

OpenStudy (luigi0210):

which simplifies to 3.. unless I've completely lost my math skills

OpenStudy (nurali):

3

OpenStudy (anonymous):

were we plug in for (-2) ^3, why is it not -6

OpenStudy (luigi0210):

cause it's -2*-2*-2

OpenStudy (luigi0210):

a^2=a*a a^3=a*a*a a^4=a*a*a*a etc..

OpenStudy (luigi0210):

and a^0=1

OpenStudy (anonymous):

oh, I see. I was doing something wrong, thanks ^^

OpenStudy (luigi0210):

Np

OpenStudy (luigi0210):

:P